OpenStudy (anonymous):

I need help on this. I'm stuck on this one because I don't know what to do with the "6". Solve the equation. Check for extraneous solutions. 6|6-4x|=8x+4

terenzreignz (terenzreignz):

Well, you basically 'get rid of it' by dividing both sides by 6 :)

OpenStudy (anonymous):

I figured so. Wasn't sure. So that leaves me with |6-4x|=8x+ 2/3?

terenzreignz (terenzreignz):

You also have to divide the 8x :)

terenzreignz (terenzreignz):

\[\Large |6-4x |= \frac43x+\frac23\]

terenzreignz (terenzreignz):

Savvy? ^_^

terenzreignz (terenzreignz):

Or, there's actually another way... You can just distribute the 6 to the absolute value...

terenzreignz (terenzreignz):

It makes no difference really, I'm just delaying dealing with fractions until there's absolutely no choice XD

terenzreignz (terenzreignz):

So, either way... \[\Large |6-4x|= \frac43x +\frac23\] or\[\Large |36-24x|= 8x + 4\]

terenzreignz (terenzreignz):

They'll work out just fine ^_^

OpenStudy (anonymous):

Thank you! Fractions do make it harder but the answers include some fractions so it would be better to do the fractions :)

terenzreignz (terenzreignz):

Fractions inevitable in this question. Can you take it from here?

OpenStudy (anonymous):

I'm working it out to see if I get stuck again. I am a little stuck though. I'm not sure what to do next

OpenStudy (anonymous):

Subtract 6 from both sides?

terenzreignz (terenzreignz):

Well, yes and no. When you have an absolute value equation, it basically gives you two equations to solve.. one which simply removes the absolute value bars... \[\Large 6-4x = \frac43x+\frac23\] and one which removes the absolute value bars BUT multiplies whatever was inside it by -1. \[\Large \color{blue}{(-1)}(6-4x )= \frac43x + \frac23\]\[\Large4x - 6 = \frac43x + \frac23\] Just solve both of these, you'll get two *possible* values for x.

OpenStudy (anonymous):

Ohh okay. I was thinking \[6-4x = \frac{ 4}{ 3 }x + \frac{2}{3} \] and \[6-4x = -(\frac{4}{3}x - \frac{2}{3}) \]

terenzreignz (terenzreignz):

No parentheses on the second. unless you have a + sign inside \[\Large 6 - 4x = -\left(\frac43x +\frac23\right)\]or\[\Large 6-4x = -\frac43x - \frac23\] But not\[\Large 6 -4x = -\left(\frac43x - \frac23\right)\color{red}{=-\frac43x + \frac23}\]

OpenStudy (anonymous):

Ohhh okay thank you. For the first one I came out with \[\frac{16}{3}x = -\frac{16}{3}\]

terenzreignz (terenzreignz):

Something's wrong here...

terenzreignz (terenzreignz):

Could you show your steps?

OpenStudy (anonymous):

\[6|6-4x|=8x+4\] Divided both sides by 6 and got \[6-4x=\frac{4}{3}x + \frac{2}{3}\] then subtracted 6 from both sides and got \[\frac{16}{3}x=-\frac{16}{3}\] then divided both sides by \[\frac{16}{3}\] and got -1.

terenzreignz (terenzreignz):

Tsk... that fateful 'wrong' sign... if in doubt, don't skip steps... patience is a virtue :D \[6-4x=\frac{4}{3}x + \frac{2}{3}\] Subtract 6 from both sides \[6-4x\color{red}{-6}=\frac{4}{3}x + \frac{2}{3}\color{red}{-6}\]\[\Large -4x = \frac43x\color{blue}{-\frac{16}3}\] Right?

OpenStudy (anonymous):

Ohhh I did skip a step!

terenzreignz (terenzreignz):

Now, if we subtract \(\Large \frac43 x\) from both sides \[\Large -4x\color{red}{-\frac43x} = \frac43x -\frac{16}3\color{red}{-\frac43x}\] We, in fact, arrive at\[\Large \color{blue}{-\frac{16}3x }=-\frac{16}3\] So... x = ? And yes, one wrong side could be tragic :D

terenzreignz (terenzreignz):

Err, one wrong sign* sorry...

OpenStudy (anonymous):

So x=1 :)

terenzreignz (terenzreignz):

Yes, and that's one of the two solutions. Now solve the other one, and for the love of ******* don't skip steps (yet) XD

OpenStudy (anonymous):

LOL I won't this time! :)